# Factoriality and birational rigidity of two families of singular quartic three-folds

### v1

@@ Introduction
0.1. The main result. The present paper is a follow up of [1]. We prove factoriality and birational rigidity and describe the group of birational automorphisms of two classes of three-dimensional quartics in ${\mathbb{P}}^{4}$ (the ground field is the field of complex numbers $\mathbb{C}$}:
@@
Let $\Sigma$ be a mobile linear system on $V$ (that is, $\mathop{\rm codim}\mathop{\rm Bs}\Sigma\geqslant 2$, the system $\Sigma$ has no fixed components). By Theorem 0.1, there is $n[\Sigma]\geqslant 1$, such that $\Sigma\subset|n[\Sigma]\cdot H|$. We say that $\Sigma$ has no maximal singularities, if the pair $\left(V,\frac{1}{n[\Sigma]}D\right)$ is canonical, that is, for every excetional divisor $E$ over $V$ the inequality
@@
The third main result of the present paper is an explicit description of the group of birational automorphisms $\mathop{\rm Bir}V$. In contrast to the first two main results, its statement is not the same for the quartics of the first and second type. This is because the second singularity of a quartiv of the first type is an infinitely near point and does not generate any additional birational automorphisms.
@@
Proposition 0.1. For a general quartic of the second type the equality $L(o_{2})\leqslant 24$ holds.
We denote the subgroup in $\mathop{\rm Bir}V$, generated by these $26+L(o_{2})$ involutions, by the symbol $B(V)$.
@@
Remark 0.2. In the semi-direct product for the group $\mathop{\rm Bir}V$ the finite subgroup $\mathop{\rm Aut}V$ acts on $B(V)$ by conjugation, preserving the point $o_{1}$ and permuting the lines, passing through the point $o_{1}$, for quartics of the first type, and preserving both points $o_{1}$, $o_{2}$ and permuting the lines passing through these points (iside each of the two groups of lines). Thus in the first family of quartics the subgroup $B(V)$ is freely generated by 25 involutions and in the second family for a general quartic it is freely generated by 50 involutions.
@@
0.5. Birational geography of three-dimensional quartics. Irreducible quartic hypersurfaces in ${\mathbb{P}}^{4}$ are parameterized by points of a Zariski open subset in the projective space ${\mathbb{P}}^{69}$, that is, by points of a quasi-projective 69-dimensional variety. In [1, Subsection 0.4] the problem of studying the ‘‘birational geography’’ of three-dimensional quartics was formulated: to break the natural parameter space into disjoint quasi-projective subvarieties, consisting of quartics with a similar birational type. The natural ‘‘measure of size’’ of these subvarieties is their dimension (or the codimension in ${\mathbb{P}}^{69}$). The seminal paper [5] that started the theory of birational rigidity described the birational type of non-singular quartics, reducing the subsequent work to investigating the 68-dimensional subset of singular quartics, that is, to the subset of codimension 1. It follows from [4] that the birational rigidity still holds in codimension 1, too (although the group of birational automorphisms becomes very large). It follows from [6] (taking into account [7, 8, 9]) and [1] that the codimension of the complement to the set of factorial birationally rigid three-dimensional quartics is at least 3. Non-rigid but ‘‘almost rigid’’ quartics were studies in [2], see also [3]. For the families and types of singularities that are already accessible for the modern techniques (however, often very hard to study), see. [1, Subsection 0.4]. Note also the paper [10], which it is natural to consider as a part of the birational geography of Fano double spaces of dimension 3, branched over a sextic surface in ${\mathbb{P}}^{3}$.
@@ 1 Factoriality
1.1. Preliminary remarks. We use the notations of Subsections 0.2, 0.3, considering in detail quartics of the first type. Note, in the first place, that the proof of factoriality, presented in [2, Remark 1.2], works for a quartics of the first type $V$, too, as by construction the homogeneous polynomial $q_{4}(z_{*})$ is arbitrary: the section of $V$ by the hypersurface ‘‘at infinity’’ is precisely $\{q_{4}=0\}$. Therefore, for a general quartic of the first type its section ‘‘at infinity’’ has Picard number 1. However, we give a different proof of factoriality that does not use this fact.
Note, first of all, that on $X_{2}$ the class of the hypersurface $V_{2}$ is $4H_{\mathbb{P}}-2{\mathbb{E}}_{1}-2{\mathbb{E}}_{2}$. This class is nef and big, but not ample, so that the Lefschetz theorem does not apply (in contrast to [1]). Furthermore, it is easy to see that all cohomology spaces
@@
and the six similar spaces for $V$, $V_{1}$, $V_{2}$ are trivial, so that (by the explonential sequences for $X_{2}$ and $V_{2}$) there are the natural isomorphisms
@@ 2 Birational rigidity
In order to prove Theorem 2.1 we have to investigate all these 8 cases. Note that essentially this theorem claims that either (1.1) or (2.1) takes place, and in the latter case only one line is the centre of a maximal singularity.
@@
Proof of the lemma. The strict transforms on $X_{1}$ of the quadrics in ${\mathbb{P}}^{4}$, cutting out the curve $B$, cut out $B^{1}$ on $X_{1}$. Therefore, $W_{2}\cap{\mathbb{E}}_{2}$ is a general plane in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, containing the point $B_{2}\cap{\mathbb{E}}_{2}$ (when we add the upper index 2, we mean taking the strict transform on $X_{2}$), so that
@@
(In fact, the computations performed above in the singular case and the compuations in [1, Subsection 2.2] in the non-singular case give a stronger estimate $\mu<\frac{1}{4}(-3+\sqrt{41})$, but we do not need it.)
@@
is a mobile linear system, however, the inetegers $m_{1}$, $m_{2}$ are not related by the ineqyality $m_{1}\geqslant m_{2}$. Let $L_{1,1},\dots,L_{1,24}$ be the lines on $V$, passing through the point $o_{1}$, and $L_{2,1},\dots,L_{2,L(o_{2})}$ the lines on $V$, passing through the point $o_{2}$. Theorem 2.1 should be replaced by a claim about $26+L(o_{2})$ multiplicities.
@@
Proof of Proposition 2.5 is put off until §4, since the point $o_{i}$ and the line $L_{j,l}$ are contained in some plane, which requires the proof of the plane section lemma.
@@
3.1. Statement of the problem. Recall that we consider quartics of the first type, so that $o_{2}$ is the vertex of the cone $Q_{1}$. Let $P\ni o_{1}$ be some two-dimensional plane, $P_{1}\subset X_{1}$ its strict transform and $R=P^{1}\cap{\mathbb{E}}_{1}$ the exceptional line of the blow up of the point $o_{1}$ on $P$. If $R$ is not a generator of the cone $Q_{1}$, then $R\not\subset Q_{1}$. Assume first that $o_{2}\not\in R$, that is, $o_{2}\not\in P^{1}$. In that case the proof of the plane section lemma in [4, §6] works without changes so that the plane section lemma is shown (see also the comments in [1, Proposition 2.1]).
@@
where $(P\circ V)$ is the plane quartic cut out by the hypersurface $V$ on the plane $P$, it can be reducible and contain multiple components. Its affine part is given by the equation $f(z_{*})|_{P}=0$. It is easy to see that $R\subset(P_{1}\circ V^{1})$ if and only if $\mu\geqslant 3$, and $R$ comes into the effective 1-cycle $(P^{1}\circ V^{1})$ with multiplicity $\mu_{2}$.
@@
Let us list all options for the cycle $(P\circ V)$. Each option has its number in the format (Pa.b), where $a=\mu$. We consider the planes, for which $(P\circ V)$ is not an irreducible curve of degree 4, such curves can not be the centre of a maximal singularity, see Subsection 2.2.
@@
Note, that $\theta_{22}=-\frac{3}{24}$, and this agrees with the fact that the line $C_{2}$ can be a maximal curve of the system $\Sigma$.
@@
Remark 3.1. The argument above, when we restrict the linear system $\Sigma$ onto a general surface $S\in|H-P|$ and study the intersection form on the verctor space, the basis of which is formed by the numerical classes of the components of the plane section and the exceptional curves of the resolution of singularities of the surface $S$, is the explicit geometric form of the argument in [4, §6], which was carried out in terms of numerical classes of a certain blow-up of the quartic. In [4, §6] it was done in that way in order to avoid considering a large number of possible cases. In this paper we restrict the amount of cases by the conditions of general position and so it is easier to prove the plane section lemma for each case, in which the point $o_{2}$ is present, by explicit geometric arguments, resolving singularities of the surface $S$ and analysing the intersection form for a particular set of curves. However, the logic of the arguments is the same as in [4, §6]; in particular, the replacement of the classes $C_{1}^{2}$, $C_{2}^{2}$, $R_{1}^{2}$ by the classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$, orthogonal to the curves $R_{2}^{2}$, $R_{3}$, corresponds to the ‘‘reductions’’ in [4, §6].
@@
(where in contrast to the previous cases $R_{1}\neq R$: the line $R\subset{\mathbb{E}}_{1}$ intersects the cone $Q_{1}$ at the only point, the vertex $o_{2}$), besides,
@@
$\left(\begin{array}[]{ccccc}-2&1&1&0&0\\ 1&-2&1&1&1\\ 1&1&-2&1&0\\ 0&1&1&-2&1\\ 0&1&0&1&-1\end{array}\right).$
@@
Lemma 4.3 In the coordinates $(z_{1},z_{2},u,z_{4})$ the quadratic part of the polynomial $f$ at the point $o_{2}$ is
@@ 5 Infinitely near lines
be the sequence of blow-ups of curves $B_{i-1}\subset V^{(i-1)}$ with the exceptional divisors $E^{(i)}\subset V^{(i)}$, $E^{(i)}=\sigma_{i}^{(-1)}(B_{i-1})$, where $i=2,\dots,k+1$, $k\geqslant 2$, such that:
@@
$m_{1}+\nu_{2}+\dots\nu_{k+1}>(k+1)n.$
Let $Z=(D_{1}\circ D_{2})$ be the self-intersection of the system $\Sigma$, where $D_{1},D_{2}\in\Sigma$ are general divisors. This is and effective 1-cycle on $V$, for which, according to the technique of counting multiplicities, see [15, Chapter 2, Proposition 2.4, Corollary 2.5], and also [15, Chapter 2, Subsection 2.2], [1, §3], we have the estimate
@@
This effective 1-cycle on $Q_{2}^{(2)}$ has multiplicity $\geqslant\nu_{2}$ at the point $p_{2}=B_{2}\cap Q_{2}^{(2)}=B_{2}\cap N$ (the inetrsections are transversal, since $B_{2}$ is a section of the ruled surface $E^{(2)}/B_{1}$).
@@
where $Y^{(2)}$ means the strict transform. This completes the proof of Proposition 5.1.
@@
(1) the case of general position, when $(V\circ P)=V\cap P$ is an irreducible curve $C$ of degree 4 with the triple point $o_{1}$, so that $o_{2}\in C^{1}$ is a non-singular point (this case takes place for allmost all generators $R$, except for finitely many of them),
@@
We will exclude these two cases in different ways: the method of excluding the case (1) does not work in the case (2), and the method of excluding the case (2) does not apply to finitely many generators $R$ and planes $P$, for which the curve $C$ has at the point $o_{1}$ two branches, a non-singular one and a cuspidal one, and moreover, the point $o_{2}$ lies exactly on the cuspidal branch. The case (1) will be excluded by means of the ‘‘quadratic’’ technique, which is more traditional and simpler. Our approach to the proof of Proposition 2.4 in the case (1) is to estimate the input of the curve $C$ to the self-intersection $Z$ of the system $\Sigma$; we will show that this input is sufficiently big, which will make it possible to obtain a contradiction by means of the standard technique of counting multiplicities:
@@
(from the inequalities (5) and (6) it follows that the right hand side is strictly positive). It is easy to check that for a fixed value $\nu_{2}+\nu_{3}=2\theta$ the estimates (7) and (8) are the worst for $\nu_{2}=\nu_{3}=\theta$ (recall that $\nu_{2}\geqslant\nu_{3}$: on a non-singular variety multiplicities do not increase), so that we may assume that this equality takes place.
@@
Write down $Z=b_{C}^{2}C+Z^{\sharp}$, where the effective 1-cycle $Z^{\sharp}$ does not contain the curve $C$ as a component. Obviously,
$\mathop{\rm deg}Z=4b_{C}^{2}+\mathop{\rm deg}Z^{\sharp}=4n^{2}$
@@
$\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant 3b_{C}^{2}+\mathop{\rm deg}Z^{\sharp}=4n^{2}-b_{C}^{2}.$
@@
and $\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant\frac{26}{9}\theta^{2}$. Again we write down $Z=b_{C}^{2}C+Z^{\sharp}$ and obtain (in the same way as above) the inequality $139\theta^{2}-300n\theta+162n^{2}\leqslant 0$, where $\theta\geqslant\frac{9}{8}n$. However the quadratic function $139t^{2}-300t+162$ is strictly increasing for $t\geqslant\frac{9}{8}$, and its value for $t=\frac{9}{8}$ is $\frac{27}{64}>0$ (up to a factor, it is the same value as above, since for $\theta=\frac{9}{8}n$ we have $\frac{2}{3}\theta=3n-2\theta$). Contradiction.
@@
where the effective 1-cycle $Y^{\sharp}$ does not contain $L_{1}$, $L_{2}$ as components and $l_{1},l_{2}\in{\mathbb{Z}}_{+}$. Obviously, $l_{1}\leqslant m_{1}$ and $l_{2}\leqslant m_{2}$ and $Y^{\sharp}$ is of bi-degree $(m_{2}-l_{1},m_{2}-l_{2})$. Now we get:
@@
$(n-b_{C})C^{@}+9n-b_{L}-L^{2}+(2n-m_{1}-\nu_{2})R^{2}_{1}+(n-m_{1})R^{2}_{2}+(2n-m_{1}-m_{2})R_{3}$
on the non-singular surface $S^{2}$. The inequalities $(\xi\cdot C^{2})\geqslant 0$ and $(\xi\cdot L^{2})\geqslant 0$ (see the multiplication table (1)) turn into the estimates
$b_{C}\geqslant\frac{1}{2}(b_{L}+2m_{1}+m_{2}+\nu_{2}-3n),b_{L}\geqslant\frac{1}{2}(b_{C}+m_{1}+\nu_{2}-n).$
@@
Список литературы
  • [1] Pukhlikov A. V., Birational rigidity of quartic three-folds with a double point of rank 3, Advances in Geometry 25 (2025), No. 4, 447–461.
  • [2] Corti A. and Mella M., Birational geometry of terminal quartic 3-folds. I. Amer. J. Math. 126 (2004), No. 4, 739–761.
  • [3] Ahmadinezhad H. and Kaloghiros A.-S., Non-rigid quartic 3-folds, Compos. Math. 152 (2016), No. 5, 955–983.
  • [4] Pukhlikov A. V., Birational automorphisms of a three-dimensional quartic with an elementary singularity, Math. USSR Sb. 63 (1989), 457–482.
  • [5] Iskovskikh V. A. and Manin Yu. I., Three-dimensional quartics and counterexamples to the Lüroth problem, Math. USSR Sb. 86 (1971), No. 1, 140–166.
  • [6] Mella M., Birational geometry of quartic 3-folds. II, Math. Ann. 330 (2004), No. 1, 107–126.
  • [7] Shramov K. A., Birational automorphisms of nodal quartic threefolds, arXiv:0803.4348, 1–32.
  • [8] Cheltsov I., Factorial threefold hypersurfaces, J. Alg. Geom. 19 (2010), 781–791.
  • [9] Shramov C.A. $\mathbb{Q}$}-factorial three-dimensional quartics, Sbornik: Mathematics 198 (2007), No. 7-8, 1165–1174.
  • [10] Krylov I., Okada T., Paemurru E. and Park J., $2n^{2}$-inequality for $cA_{1}$ points and applications to birational rigidity, Compos. Math. 160 (2024), No. 7, 1551–1595.
  • [11] Akizuki Y. and Nakano S., Note on Kodaira-Spencer’s proof of Lefschetz’s theorem, Proc. Jap. Acad., Ser. A 30 (1954), 266–272.
  • [12] Esnault H. and Viehweg E., Lectures on vanishing theorems, DMV Seminar, Band 20. Birkhäuser Verlag, Basel-Bpstpn-Berlin. 1992. 164 pp.
  • [13] Corti A., Singularities of linear systems and 3-fold birational geometry, in ‘‘Explicit Birational Geometry of Threefolds’’, London Mathematical Society Lecture Note Series 281 (2000), Cambridge University Press, 259–312.
  • [14] Pukhlikov A. V., Essentials of the method of maximal singularities, in ‘‘Explicit Birational Geometry of Threefolds’’, London Math. Soc. Lecture Note Ser. 281 (2000), Cambridge University Press, 73–100.
  • [15] Pukhlikov A., Birationally Rigid Varieties. Mathematical Surveys and Monographs 190, AMS, Providence, RI, 2013, 365 pp.
Department of Mathematical Sciences,
The University of Liverpool
pukh@liverpool.ac.uk

### v2

@@ Introduction
0.1. The main result. The present paper is a follow up of [1]. We prove factoriality and birational rigidity and describe the group of birational automorphisms of two classes of three-dimensional quartics in ${\mathbb{P}}^{4}$ (the ground field is the field of complex numbers $\mathbb{C}$}):
@@
Let $\Sigma$ be a mobile linear system on $V$ (that is, $\mathop{\rm codim}\mathop{\rm Bs}\Sigma\geqslant 2$, the system $\Sigma$ has no fixed components). By Theorem 0.1, there is $n[\Sigma]\geqslant 1$, such that $\Sigma\subset|n[\Sigma]\cdot H|$. We say that $\Sigma$ has no maximal singularities, if the pair $\left(V,\frac{1}{n[\Sigma]}D\right)$ is canonical, that is, for every exceptional divisor $E$ over $V$ the inequality
@@
The third main result of the present paper is an explicit description of the group of birational automorphisms $\mathop{\rm Bir}V$. In contrast to the first two main results, its statement is not the same for the quartics of the first and second type. This is because the second singularity of a quartic of the first type is an infinitely near point and does not generate any additional birational automorphisms.
@@
Proposition 0.1. For a general quartic of the second type the equality $L(o_{2})=24$ holds.
We denote by the symbol $B(V)$ the subgroup in $\mathop{\rm Bir}V$, generated by these $26+L(o_{2})$ involutions.
@@
Remark 0.2. In the semi-direct product for the group $\mathop{\rm Bir}V$ the finite subgroup $\mathop{\rm Aut}V$ acts on $B(V)$ by conjugation, preserving the point $o_{1}$ and permuting the lines, passing through the point $o_{1}$, for quartics of the first type, and preserving both points $o_{1}$, $o_{2}$ and permuting the lines passing through these points (inside each of the two groups of lines). Thus in the first family of quartics the subgroup $B(V)$ is freely generated by 25 involutions and in the second family for a general quartic it is freely generated by 50 involutions.
@@
0.5. Birational geography of three-dimensional quartics. Irreducible quartic hypersurfaces in ${\mathbb{P}}^{4}$ are parameterized by points of a Zariski open subset in the projective space ${\mathbb{P}}^{69}$, that is, by points of a quasi-projective 69-dimensional variety. In [1, Subsection 0.4] the problem of studying the “birational geography” of three-dimensional quartics was formulated: to break the natural parameter space into disjoint quasi-projective subvarieties, consisting of quartics with a similar birational type. The natural “measure of size” of these subvarieties is their dimension (or the codimension in ${\mathbb{P}}^{69}$). The seminal paper [5] that started the theory of birational rigidity described the birational type of non-singular quartics, reducing the subsequent work to investigating the 68-dimensional subset of singular quartics, that is, to the subset of codimension 1. It follows from [4] that the birational rigidity still holds in codimension 1, too (although the group of birational automorphisms becomes very large). It follows from [6] (taking into account [7, 8, 9]) and [1] that the codimension of the complement to the set of factorial birationally rigid three-dimensional quartics is at least 3. Non-rigid but “almost rigid” quartics were studies in [2], see also [3]. For the families and types of singularities that are already accessible for the modern techniques (however, often very hard to study), see. [1, Subsection 0.4]. Note also the paper [10], which it is natural to consider as a part of the birational geography of Fano double spaces of dimension 3, branched over a sextic surface in ${\mathbb{P}}^{3}$.
@@ 1 Factoriality
1.1. Preliminary remarks. We use the notations of Subsections 0.2, 0.3, considering in detail quartics of the first type. Note, in the first place, that the proof of factoriality, presented in [2, Remark 1.2], works for a quartics of the first type $V$, too, as by construction the homogeneous polynomial $q_{4}(z_{*})$ is arbitrary: the section of $V$ by the hypersurface “at infinity” is precisely $\{q_{4}=0\}$. Therefore, for a general quartic of the first type its section “at infinity” has Picard number 1. However, we give a different proof of factoriality that does not use this fact.
Note, first of all, that on $X_{2}$ the class of the hypersurface $V_{2}$ is $4H_{\mathbb{P}}-2{\mathbb{E}}_{1}-2{\mathbb{E}}_{2}$. This class is nef and big, but not ample, so that the Lefschetz theorem does not apply (in contrast to the quartics considered in [1]). Furthermore, it is easy to see that all cohomology spaces
@@
and the six similar spaces for $V$, $V_{1}$, $V_{2}$ are trivial, so that (by the exponential sequences for $X_{2}$ and $V_{2}$) there are the natural isomorphisms
@@ 2 Birational rigidity
In order to prove Theorem 2.1, we have to investigate all these 8 cases. Note that essentially this theorem claims that either (1.1) or (2.1) takes place, and in the latter case only one line is the centre of a maximal singularity.
@@
Proof of the lemma. The strict transforms on $X_{1}$ of the quadrics in ${\mathbb{P}}^{4}$, cutting out the curve $B$, cut out $B^{1}$ on $X_{1}$. Therefore, $W_{2}\cap{\mathbb{E}}_{2}$ is a general plane in ${\mathbb{E}}_{2}\cong{\mathbb{P}}^{3}$, containing the point $B^{2}\cap{\mathbb{E}}_{2}$ (when we add the upper index 2, we mean taking the strict transform on $X_{2}$), so that
@@
(In fact, the computations performed above in the singular case and the computations in [1, Subsection 2.2] in the non-singular case give a stronger estimate $\mu<\frac{1}{4}(-3+\sqrt{41})$, but we do not need it.)
@@
is a mobile linear system, however, the integers $m_{1}$, $m_{2}$ are not related by the inequality $m_{1}\geqslant m_{2}$. Let $L_{1,1},\dots,L_{1,24}$ be the lines on $V$, passing through the point $o_{1}$, and $L_{2,1},\dots,L_{2,L(o_{2})}$ the lines on $V$, passing through the point $o_{2}$. Theorem 2.1 should be replaced by a claim about $26+L(o_{2})$ multiplicities.
@@
Proof of Proposition 2.5 we put off until §4, since the point $o_{i}$ and the line $L_{j,l}$ are contained in some plane, which requires the proof of the plane section lemma.
@@
3.1. Statement of the problem. Recall that we consider quartics of the first type, so that $o_{2}$ is the vertex of the cone $Q_{1}$. Let $P\ni o_{1}$ be some two-dimensional plane, $P^{1}\subset X_{1}$ its strict transform and $R=P^{1}\cap{\mathbb{E}}_{1}$ the exceptional line of the blow up of the point $o_{1}$ on $P$. If $R$ is not a generator of the cone $Q_{1}$, then $R\not\subset Q_{1}$. Assume first that $o_{2}\not\in R$, that is, $o_{2}\not\in P^{1}$. In that case the proof of the plane section lemma in [4, §6] works without changes so that the plane section lemma is shown (see also the comments in [1, Proposition 2.1]).
@@
where $(P\circ V)$ is the plane quartic cut out by the hypersurface $V$ on the plane $P$, it can be reducible and contain multiple components. Its affine part is given by the equation $f(z_{*})|_{P}=0$. It is easy to see that $R\subset(P^{1}\circ V^{1})$ if and only if $\mu\geqslant 3$, and $R$ comes into the effective 1-cycle $(P^{1}\circ V^{1})$ with multiplicity $\mu-2$.
@@
Let us list all options for the cycle $(P\circ V)$. Each option has its number in the format (a.b), where $a=\mu$. We consider the planes, for which $(P\circ V)$ is not an irreducible curve of degree 4, such curves can not be the centre of a maximal singularity, see Subsection 2.2.
@@
Note that $\theta_{22}=-\frac{3}{24}$, and this agrees with the fact that the line $C_{2}$ can be a maximal curve of the system $\Sigma$.
@@
Remark 3.1. The argument above, when we restrict the linear system $\Sigma$ onto a general surface $S\in|H-P|$ and study the intersection form on the vector space, the basis of which is formed by the numerical classes of the components of the plane section and the exceptional curves of the resolution of singularities of the surface $S$, is the explicit geometric form of the argument in [4, §6], which was carried out in terms of numerical classes of a certain blow-up of the quartic. In [4, §6] it was done in that way in order to avoid considering a large number of possible cases. In this paper we restrict the amount of cases by the conditions of general position and so it is easier to prove the plane section lemma for each case, in which the point $o_{2}$ is present, by explicit geometric arguments, resolving singularities of the surface $S$ and studying the intersection form for a particular set of curves. However, the logic of the arguments is the same as in [4, §6]; in particular, the replacement of the classes $C_{1}^{2}$, $C_{2}^{2}$, $R_{1}^{2}$ by the classes $C_{1}^{+}$, $C_{2}^{+}$, $R_{1}^{+}$, orthogonal to the curves $R_{2}^{2}$, $R_{3}$, corresponds to the “reductions” in [4, §6].
@@
(where, in contrast to the previous cases, $R_{1}\neq R$: the line $R\subset{\mathbb{E}}_{1}$ intersects the cone $Q_{1}$ at the only point, the vertex $o_{2}$), besides,
@@
$\left(\begin{array}[]{ccccc}-2&1&1&0&0\\ 1&-2&1&1&1\\ 1&1&-2&1&0\\ 0&1&1&-2&1\\ 0&1&0&1&-2\end{array}\right).$
@@
Lemma 4.3. In the coordinates $(z_{1},z_{2},u,z_{4})$ the quadratic part of the polynomial $f$ at the point $o_{2}$ is
@@ 5 Infinitely near lines
be the sequence of blow-ups of curves $B_{i-1}\subset V^{(i-1)}$ with the exceptional divisors $E^{(i)}\subset V^{(i)}$, $E^{(i)}=\sigma_{i}^{-1}(B_{i-1})$, where $i=2,\dots,k+1$, $k\geqslant 2$, such that:
@@
$m_{1}+\nu_{2}+\dots+\nu_{k+1}>(k+1)n.$
Let $Z=(D_{1}\circ D_{2})$ be the self-intersection of the system $\Sigma$, where $D_{1},D_{2}\in\Sigma$ are general divisors. This is and effective 1-cycle on $V$, for which, according to the technique of counting multiplicities, see [15, Chapter 2, Proposition 2.4, Corollary 2.5], and the computation at the end of [15, Chapter 2, Subsection 2.2], and also [1, §3], we have the estimate
@@
This effective 1-cycle on $Q_{2}^{(2)}$ has multiplicity $\geqslant\nu_{3}$ at the point $p_{2}=B_{2}\cap Q_{2}^{(2)}=B_{2}\cap N$ (the intersections are transversal, since $B_{2}$ is a section of the ruled surface $E^{(2)}/B_{1}$).
@@
where $Y^{(2)}$ means the corresponding strict transform. This completes the proof of Proposition 5.1.
@@
(1) the case of general position, when $(V\circ P)=V\cap P$ is an irreducible curve $C$ of degree 4 with the triple point $o_{1}$, so that $o_{2}\in C^{1}$ is a non-singular point (this case takes place for almost all generators $R$, except for finitely many of them),
@@
We will exclude these two cases in different ways: the method of excluding the case (1) does not work in the case (2), and the method of excluding the case (2) does not apply to finitely many generators $R$ and planes $P$, for which the curve $C$ has at the point $o_{1}$ two branches, a non-singular one and a cuspidal one, and moreover, the point $o_{2}$ lies exactly on the cuspidal branch. The case (1) will be excluded by means of the “quadratic” technique, which is more traditional and simpler. Our approach to the proof of Proposition 2.4 in the case (1) is to estimate the input of the curve $C$ to the self-intersection $Z$ of the system $\Sigma$; we will show that this input is sufficiently big, which will make it possible to obtain a contradiction by means of the standard technique of counting multiplicities:
@@
(It follows from the inequalities (5) and (6) that the right hand side is strictly positive.) It is easy to check that for a fixed value $\nu_{2}+\nu_{3}=2\theta$ the estimates (7) and (8) are the worst for $\nu_{2}=\nu_{3}=\theta$ (recall that $\nu_{2}\geqslant\nu_{3}$: on a non-singular variety multiplicities do not increase), so that we may assume that this equality takes place.
@@
Write down $Z=\beta C+Z^{\sharp}$, where the effective 1-cycle $Z^{\sharp}$ does not contain the curve $C$ as a component and $\beta\geqslant b_{C}^{2}$. Obviously,
$\mathop{\rm deg}Z=4\beta+\mathop{\rm deg}Z^{\sharp}=4n^{2}$
@@
$\mathop{\rm mult}\nolimits_{o_{1}}Z\leqslant 3\beta+\mathop{\rm deg}Z^{\sharp}=4n^{2}-\beta.$
@@
and $\mathop{\rm mult}\nolimits_{o_{1}}Z\geqslant\frac{26}{9}\theta^{2}$. Again we write down $Z=\beta C+Z^{\sharp}$ with $\beta\geqslant b_{C}^{2}$ and obtain (in the same way as above) the inequality $139\theta^{2}-300n\theta+162n^{2}\leqslant 0$, where $\theta\geqslant\frac{9}{8}n$. However the quadratic function $139t^{2}-300t+162$ is strictly increasing for $t\geqslant\frac{9}{8}$, and its value for $t=\frac{9}{8}$ is $\frac{27}{64}>0$ (up to a factor, it is the same value as above, since for $\theta=\frac{9}{8}n$ we have $\frac{2}{3}\theta=3n-2\theta$). Contradiction.
@@
where the effective 1-cycle $Y^{\sharp}$ does not contain $L_{1}$, $L_{2}$ as components and $l_{1},l_{2}\in{\mathbb{Z}}_{+}$. Obviously, $l_{1}\leqslant m_{2}$ and $l_{2}\leqslant m_{2}$ and $Y^{\sharp}$ is of bi-degree $(m_{2}-l_{1},m_{2}-l_{2})$. Now we get:
@@
$(n-b_{C})C^{2}+(n-b_{L})L^{2}+(2n-m_{1}-\nu_{2})R^{2}_{1}+(n-m_{1})R^{2}_{2}+(2n-m_{1}-m_{2})R_{3}$
on the non-singular surface $S^{2}$. The inequalities $(\xi\cdot C^{2})\geqslant 0$ and $(\xi\cdot L^{2})\geqslant 0$ turn into the estimates (see the multiplication table (1))
$b_{C}\geqslant\frac{1}{2}(b_{L}+2m_{1}+m_{2}+\nu_{2}-3n),\quad b_{L}\geqslant\frac{1}{2}(b_{C}+m_{1}+\nu_{2}-n).$